solve using system of elimination.
4x-3y=11
3x-5y+-11
wut r the steps to solve this equation

Answer by bacos(9)

Simplifying
4x + -3y = 11
Solving
4x + -3y = 11
Solving for variable 'x'.
Move all terms containing x to the left, all other terms to the right.
Add '3y' to each side of the equation.
4x + -3y + 3y = 11 + 3y
Combine like terms: -3y + 3y = 0
4x + 0 = 11 + 3y
4x = 11 + 3y
Divide each side by '4'.
x = 2.75 + 0.75y
Simplifying
x = 2.75 + 0.75y
Simplifying
3x + -5y + -11
Reorder the terms:
-11 + 3x + -5y